OpenStudy (felisha):

Superman needs to save Lois from the clutches of Lex Luthor. After flying for 5 seconds, he is 1900 meters from her. Then at 17 seconds he is 1180 meters from her. Write an equation to model this situation (use m for meters and s for seconds)

OpenStudy (erak):

Just wondering, do you know what kind of function it is? e.g. linear vs quadratic

OpenStudy (felisha):

Linear

OpenStudy (erak):

Okay so you would do the slope equation. Do you know what that is and how to get slope from your two points (5, 1900) and (17, 1180)?

OpenStudy (felisha):

y2-y1/x2-x1 and then y-y1=m(x-x1)

OpenStudy (erak):

So yeah, just find your m = y2-y1/x2-x1 first, and tell me what you get.

OpenStudy (felisha):

-720/12

OpenStudy (erak):

Oh sorry, didn't notice. So yeah, if you divide that, you will get -60 as your slope. Right now your equation is: y = -60x + b Substitute one of your points to solve for b.

OpenStudy (felisha):

-720/12 simplifies to -60 which becomes the slope y-1900=-60(x-5) distribute and then it is y-1900=-60x+300 add 1900 to both sides and it becomes y=-60x+2200 is that correct

OpenStudy (erak):

yes that's correct

OpenStudy (felisha):

Thank you for helping me